OpenStudy (anonymous):

1. Write the equation of the line parallel to the line 10x - 5y = 8 and passing through the point (2,4). 2. Write the equation of the line parallel to the line 4x + 2y = 5 and passing through the point (-3,5). 3. Write the equation of the line parallel to the line 4x + y = -1 and passing through the point (5,0). 4. Write the equation of the line perpendicular to the line x - 5y = -10 and passing through the point (2,5). 5. Write the equation of the line perpendicular to the line x - 3y = 9 and passing through the point (3,5). a. y = -5x + 15 b. y = 2x c. y = -3x + 14 d. y = -2x

OpenStudy (anonymous):

this is matching top with bottom

OpenStudy (texaschic101):

1. 10x - 5y = 8 -5y = -10x + 8 y = (-10x + 8) / -5 y = 2x - 8/5 The slope of this equation is 2. A parallel line will have the same slope...2 Now use the point slope formula: y - y1 = m(x - x1) with m = 2, and (2,4) x1 = 2 and y1 = 4 sub it in y - 4 = 2(x - 2) y - 4 = 2x - 4 y = 2x - 4 + 4 y = 2x (B) =============== do you understand how I did that ?

OpenStudy (texaschic101):

# 2 and 3 are done the same way. Put your equation in y = mx + b form, where m is your slope.Then sub the slope you found and your points into the point slope formula, and then solve for y.

OpenStudy (texaschic101):

# 4 and 5 are a little different because you are looking for a perpendicular line. You start out the same and put the equation in y = mx + b form. The slope you find will then have to be turned into the negative reciprocal. For instance, if you have a slope of 3/2, you take the negative reciprocal (flip the fraction and change the sign). The negative reciprocal of 3/2 is -2/3. You take the negative reciprocal and your points and sub it into the point sope formula and solve for y.

OpenStudy (anonymous):

yes i understand how to do all that just this is a matching problem

OpenStudy (anonymous):

oh never mind i understand what you did thank you

OpenStudy (texaschic101):

your welcome.....I answered your last question , just letting you know :)

OpenStudy (anonymous):

so 1 is B 2 is e 3 is d 4 is a 5 is c am i right?

OpenStudy (anonymous):

a. y = -5x + 15 b. y = 2x c. y = -3x + 14 d. y = -2x - 1 e. y = -4x + 20 this is the items to match

OpenStudy (texaschic101):

check 2 and 3 again :)